# Logarithmic Sobolev inequality in manifolds with nonnegative curvature via the ABP method

### v1

@@ Abstract.
In this paper, we employ the ABP method developed by Brendle [5] to establish the optimal $L^{p}$-logarithmic-Sobolev inequality on manifolds with nonnegative Ricci curvature, as well as a sharp $L^{2}$-logarithmic-Sobolev inequality for submanifolds in manifolds with nonnegative sectional curvature. The sharp constants in both inequalities depend on the asymptotic volume ratio of the ambient manifold.
@@ 1. Introduction
The classical sharp logarithmic Sobolev inequality on $\mathbb{R}^{n}$, proved by Gross [14], establishes that
@@
where $d\gamma=(2\pi)^{-\frac{n}{2}}e^{-\frac{|x|^{2}}{2}}\,dx$ is the standard Gaussian probability measure and $f\in C_{c}^{\infty}(\mathbb{R}^{n})$. This is equivalent to the optimal $L^{2}$-logarithmic Sobolev inequality established by Weissler [20]:
@@
The following Euclidean $L^{p}$-logarithmic Sobolev inequality, established by Beckner [2] for $p=1$, Del Pino and Dolbeault [11] for $1<p<n$ and Gentil [13] for general $p>1$, asserts that for $f\in C_{c}^{\infty}(\mathbb{R}^{n})$ with $\int_{\mathbb{R}^{n}}f^{p}\,dx=1$,
@@
Recently, by generalizing the optimal transport approach developed by Cordero-Erausquin, Nazaret and Villani [10], Kristály [16] obtained the optimal $L^{p}$-logarithmic Sobolev inequality for manifolds with nonnegative Ricci curvature. In this paper, inspired by the ABP method developed by Brendle [5], our first goal is to provide a ABP-based proof of this inequality.
@@
The sharp $L^{1}$-logarithmic Sobolev inequality for manifolds with nonnegative Ricci curvature, proved using the ABP method developed by Brendle [5], reads as follows:
@@
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with boundary $\partial\Omega$, and let $f$ be a positive smooth function in $\Omega\setminus\partial\Omega$. Then
@@
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with boundary $\partial\Omega$, and let $f$ be a positive smooth function in $\Omega\setminus\partial\Omega$ that vanishes on $\partial\Omega$. For $p>1$, we have
@@
We now turn to the logarithmic Sobolev inequality for submanifolds. In [12], Ecker derived such an inequality valid for submanifolds in Euclidean space. Later in 2020, Brendle [4] established a sharp logarithmic Sobolev inequality for compact submanifolds in Euclidean space without boundary, employing the ABP techniques developed in [3]. In 2021, Yi and Zheng [22] proved a logarithmic Sobolev inequality for compact boundaryless submanifolds in manifolds with nonnegative sectional curvature. Recently, by developing a noncompact modification of Brendle’s ABP method, Wang, Xia and Zhang [18] obtained an optimal logarithmic Sobolev inequality for non-compact self-shrinkers in Euclidean space. Independently, Balogh and Kristály [1] established the sharp logarithmic Sobolev inequality for submanifolds in Euclidean space using an optimal transport approach. Inspired by [4, 5, 22], our second goal herein is to provide an ABP-based proof of the optimal logarithmic Sobolev inequality for submanifolds in manifolds with nonnegative sectional curvature.
@@
$\int_{\Sigma}f^{2}\log f^{2}\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\theta^{-\frac{n}{2}}\int_{\Sigma}|D^{\Sigma}f|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\Big).$
@@ 2. The case of domains
without loss of generality, we assume hereafter that $\Omega$ is connected. Let $\alpha>0$ be a given parameter and $r_{o}=d(o,\Omega)$. By sacling, we may assume that
@@
$\int_{\Omega}f\log f\,d\text{vol}\leq-n-\log(\theta\omega_{n}\alpha^{-n})+\frac{1}{\alpha}\Big(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\Big)$
@@
$\int_{\Omega}f^{p}\bigg(\log f^{p}+n+\log\theta-\frac{n}{q}\log\Big(\frac{p}{q}\alpha^{\frac{q}{p}}\Big)+\log\Big(\omega_{n}\Gamma\Big(\frac{n}{q}+1\Big)\Big)\bigg)\,d\mathrm{vol}$
@@ 3. The case of submanifolds
It suffies to prove Theorem 1.3 for the case where $\Sigma$ is connected. Let $\alpha>0$ be a given parameter and $r_{o}=d(o,\Sigma)$. By scaling, we may assume that
@@
By combining Lemma 3.1 and Lemma 3.3 with the formula for change of variables in multiple integrals, we obtain
@@
Combining this inequality with 4, one can get
$\int_{\Sigma}f^{2}\Big(\log f^{2}+n+\frac{n}{2}\log\frac{\pi}{\alpha}+\log\theta\Big)\,d\mathrm{vol}-\frac{1}{\alpha}\int_{\Sigma}|D^{\Sigma}f|^{2}\,d\mathrm{vol}$
@@
$\int_{\Sigma}f^{2}\log f^{2}\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\theta^{-\frac{n}{2}}\int_{\Sigma}|Df|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\Big).$

### v2

@@ Abstract.
In this paper, we employ the ABP method developed by Brendle [5] to establish the optimal $L^{p}$ logarithmic Sobolev inequality on manifolds with nonnegative Ricci curvature, as well as a sharp $L^{2}$ logarithmic Sobolev inequality for submanifolds in manifolds with nonnegative sectional curvature. The sharp constants in both inequalities depend on the asymptotic volume ratio of the ambient manifold.
@@ 1. Introduction
The classical logarithmic Sobolev inequality on $\mathbb{R}^{n}$, which was established by Stam [19] and Federbush [13], and later highlighted by Gross [15] through its equivalence to the hypercontractivity of a diffusion semigroup, is stated as:
@@
where $d\gamma=(2\pi)^{-\frac{n}{2}}e^{-\frac{|x|^{2}}{2}}\,dx$ is the standard Gaussian probability measure and $f\in C_{c}^{\infty}(\mathbb{R}^{n})$. This is equivalent to the optimal $L^{2}$ logarithmic Sobolev inequality established by Weissler [22]:
@@
The following Euclidean $L^{p}$ logarithmic Sobolev inequality, established by Beckner [2] for $p=1$, Del Pino and Dolbeault [11] for $1<p<n$ and Gentil [14] for general $p>1$, asserts that for $f\in C_{c}^{\infty}(\mathbb{R}^{n})$ with $\int_{\mathbb{R}^{n}}f^{p}\,dx=1$,
@@
Recently, by generalizing the optimal transport approach developed by Cordero-Erausquin, Nazaret and Villani [10], Kristály [17] obtained the optimal $L^{p}$ logarithmic Sobolev inequality for manifolds with nonnegative Ricci curvature. In this paper, inspired by the ABP method developed by Brendle [5], our first goal is to provide a ABP-based proof of this inequality.
@@
The sharp $L^{1}$ logarithmic Sobolev inequality for manifolds with nonnegative Ricci curvature, proved using the ABP method developed by Brendle [5], reads as follows:
@@
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with smooth boundary $\partial\Omega$, and let $f$ be a positive smooth function in $\Omega\setminus\partial\Omega$. Then
@@
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with smooth boundary $\partial\Omega$, and let $f$ be a positive smooth function in $\Omega\setminus\partial\Omega$ that vanishes on $\partial\Omega$. For $p>1$, we have
@@
We now turn to the logarithmic Sobolev inequality for submanifolds. In [12], Ecker derived such an inequality valid for submanifolds in Euclidean space. Later in 2020, Brendle [4] established a sharp logarithmic Sobolev inequality for compact submanifolds in Euclidean space without boundary, employing the ABP techniques developed in [3]. In 2021, Yi and Zheng [24] proved a logarithmic Sobolev inequality for compact boundaryless submanifolds in manifolds with nonnegative sectional curvature. Recently, by developing a noncompact modification of Brendle’s ABP method, Wang, Xia and Zhang [20] obtained the optimal logarithmic Sobolev inequality for non-compact self-shrinkers in Euclidean space. Independently, Balogh and Kristály [1] established the sharp logarithmic Sobolev inequality for submanifolds in Euclidean space using an optimal transport approach. Inspired by [4, 5, 24], our second goal herein is to provide an ABP-based proof of the optimal logarithmic Sobolev inequality for submanifolds in manifolds with nonnegative sectional curvature.
@@
$\int_{\Sigma}f^{2}\log f^{2}\,d\mathrm{vol}\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\theta^{-\frac{2}{n}}\int_{\Sigma}|D^{\Sigma}f|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\Big).$
@@ 2. The case of domains
without loss of generality, we assume hereafter that $\Omega$ is connected. Let $\alpha>0$ be a given parameter and $r_{o}=d(o,\Omega)$. By scaling, we may assume that
@@
$\int_{\Omega}f\log f\,d\text{vol}\leq-n-\log(\theta\omega_{n}\alpha^{-n})+\frac{1}{\alpha}\Big(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\Big).$
@@
(4) $\int_{\Omega}f^{p}\bigg(\log f^{p}+n+\log\theta-\frac{n}{q}\log\Big(\frac{p}{q}\alpha^{\frac{q}{p}}\Big)+\log\Big(\omega_{n}\Gamma\Big(\frac{n}{q}+1\Big)\Big)\bigg)\,d\mathrm{vol}$
@@ 3. The case of submanifolds
It suffices to prove Theorem 1.3 for the case where $\Sigma$ is connected. Let $\alpha>0$ be a given parameter and $r_{o}=d(o,\Sigma)$. By scaling, we may assume that
@@
Combining Lemma 3.1 and Lemma 3.3 with the formula for change of variables in multiple integrals, we obtain
@@
Combining this inequality with (5), one can get
(7) $\int_{\Sigma}f^{2}\Big(\log f^{2}+n+\frac{n}{2}\log\frac{\pi}{\alpha}+\log\theta\Big)\,d\mathrm{vol}-\frac{1}{\alpha}\int_{\Sigma}|D^{\Sigma}f|^{2}\,d\mathrm{vol}$
@@
$\int_{\Sigma}f^{2}\log f^{2}\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\theta^{-\frac{2}{n}}\int_{\Sigma}|Df|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\Big).$
Acknowledgements
We would like to thank Professor Frank Morgan for his useful correspondence.